Having issues trying to solve, any help is greatly appreciated.
Given the function f(x) = -7x^2 + 14x + 4, if x = 1, what is f(x)? What point is on the graph of f?

Answer by rothauserc(4718)

f(x) = -7x^2 + 14x + 4
if x = 1, then
f(1) = -7(1^2) +14(1) + 4
f(1) = -7 + 14 +4 = 11
point (1, 11) is on the graph of f